Consider a wind energy project that produces 594 GWh/yr and is able to sell this at a cost of $0.12/kWh. The project will also receive income in this year from the PTC, at a value of $0.024/kWh. The annual operating expenses come out to be $0.009/kWh plus a fixed amount of $216 thousand. The project is able to depreciate 11.52% of its total installed cost (which was $356 million) in this year, and it can write off the interest paid on the loan, which has a value of $14.8 million this year. The principal paid on debt is $19 million and the tax rate for the company which owns the project is 35%. What is the total taxable income (TI) in millions of dollars? Round answer to two decimal places (answer may be negative).
Answer #1

Answer:

Total taxable income for the year is

85.54-5.56-55.81= 24.17  million dollars

Working:

Total revenue for this is

594000000kwh*(0.024+0.12)=85.54 dollars

Total operating expenses is

594000000kwh*0.009+216000=5.56 Millon dollars

Depreciation is 11.52% of total installed cost.

356 millions*11.52%=41.01 million dollars

Total of depreciation and interest for this year is

41.01 +14.8 =55.81 million